OpenStudy (anonymous):

Which property is illustrated in this problem? 33 + 0 = 33

OpenStudy (midhun.madhu1987):

Additive identity

OpenStudy (midhun.madhu1987):

a+0=0+a=a

OpenStudy (andrewthecookie):

"zambie" LOL